write a balanced equation for the reaction of the active ingredient in Tums with excess acid.
1
Expert's answer
2015-10-29T07:23:08-0400
There are two active components in Tums: magnesium carbonate and calcium carbonate. Normally, they are used to neutralize the main component of gastric acid - hydrochloric acid:
MgCO3 + 2HCl = MgCl2 + H2O + CO2;
CaCO3 + 2HCl = CaCl2 + H2O + CO2